# Learning Unlabeled Clients Divergence via Anchor Model Aggregation for Federated Semi-supervised Learning

### v1

@@ 1 Introduction
Federated learning has emerged as a promising solution for learning in decentralized environments, where data centralization is often infeasible due to privacy concerns. Federated learning has gained considerable attention in machine learning tasks in natural image domains (McMahan et al., 2017) as well as medical image domains (Liu et al., 2021; Saha et al., 2023; Jiang et al., 2023). Due to the challenges of data and label heterogeneity, multiple methods such as MOON (Li et al., 2021a), FedDisco (Ye et al., 2023), FedFed (Yang et al., 2023), and FedRoD (Chen & Chao, 2022) have been developed, utilizing FedAvg (McMahan et al., 2017) as their baseline. While these methods have shown promise, they often assume that all clients have exhaustive and expert-level annotations, which is a requirement that is both time-consuming and labor-intensive. This renders them impractical in real-world cross-silo federated settings, such as those found in hospitals. Therefore, developing methods that require minimal expert-level annotations in decentralized settings is a crucial area of research that deserves further attention.
Federated Semi-Supervised Learning (FedSemi) arises to alleviate exhaustive expert-level annotations by utilizing unlabeled data along with labeled data to benefit the global performance in many recent works (Zhang et al., 2023b; Cho et al., 2023; Zhang et al., 2023a; Wang et al., 2020; Jeong et al., 2021; Lin et al., 2021; Liu et al., 2021; Saha et al., 2023; Liang et al., 2022; Li et al., 2023; Kim et al., 2022; 2023b). Researchers have explored various settings for FedSemi, with the most generalizable setting considering fully unlabeled clients and other clients that can be either labeled or partially labeled, in either independent and identically distributed (IID) or non-IID settings (Liang et al., 2022; Li et al., 2023).
@@
Despite the effectiveness of these methods, existing approaches have two crucial limitations. First, their aggregation typically follows the standard FedAvg (McMahan et al., 2017), which fails to account for the fact that labeled clients can generate more accurate information than unlabeled clients. For example, when the number of unlabeled clients is dominant, the global model becomes heavily influenced by the unlabeled clients, leading to limited performance primarily due to errors introduced by the pseudo-labels assigned to the unlabeled data. Second, existing methods treat the divergence of unlabeled clients, calculated in the gradient space, as noise due to the unreliability of the pseudo-labels used in training. It is worth mentioning that client divergence may reflect the presence of minority classes and unique attributes within the dataset, rather than mere noise.
@@
To this end, we propose a novel aggregation, SemiAnAgg, which learns the importance of unlabeled clients via a consistently randomly initialized model. Our key idea is that by feeding the local client data to the same global model and the same consistently initialized random anchor model, we can measure the importance of each unlabeled client. Specifically, our intuition is two-fold: 1) The optimal feature representation for client data should be distant from random. Therefore, clients contributing to a more distant step are more likely to the global optima. 2) Since all clients share the same global and anchor model, the inter-client data variability can be measured consistently. In other words, the expectation of the global model on client data that is more distant indicates a more diverse distribution covering unique information. Consequently, SemiAnAgg gives higher weights to these clients, harnessing their potential. Without requiring any direct sharing of data, features, or gradient information among the clients and maintaining the same communication costs as previous FedSemi approaches with minimal computation and storage overhead, SemiAnAgg achieves state-of-the-art results on four widely used benchmarks with improvements up to 9% in accuracy on CIFAR-100, 9.5% on the imbalanced CIFAR-100LT, and a 7.65% improvement in recall on the medical dataset ISIC-18.
@@ 2.2 Federated Semi-Supervised Learning
Federated Semi-Supervised Learning (FedSemi) addresses the decentralized learning of both unlabeled and labeled data while preserving privacy. FedSemi has been explored in three different settings. In the first two settings, labeled data is available on the server, as in FedMatch (Jeong et al., 2021) and FedLID (Psaltis et al., 2023), or clients have partially labeled data, as seen in SemiFed (Lin et al., 2021) and FedLabel (Cho et al., 2023). The third setting, which we consider more realistic, involves fully unlabeled clients, while other clients can be either fully labeled or partially labeled. For the third setting, recent works have shifted from addressing clients with independent and identically distributed (IID) data, such as FedConsist (Wang et al., 2020) and FedIRM (Liu et al., 2021), to non-IID data, including RSCFed (Liang et al., 2022) and CBAFed (Li et al., 2023). This work focuses on the third setting, characterized by most clients with fully unlabeled and potentially non-IID data.
@@ 2.3 Federated Model Aggregation
Existing FedSemi (Liang et al., 2022; Li et al., 2023) methods regard non-robust clients or data as outliers and opt to minimize their influence. RSCFed (Liang et al., 2022) proposes to leverage unlabeled gradient divergence to eliminate non-robust (noisy) clients through average consensus inspired by RANSAC (Fischler & Bolles, 1981). However, RSCFed (Liang et al., 2022) does not consider cases when unlabeled clients are diverse due to unique information (i.e. underrepresented classes or attributes). CBAFed (Li et al., 2023) introduced federated class adaptive pseudo labeling, which can be seen as curriculum-paced pseudo learning from a global perspective (Zhang et al., 2021). It uses hard temporal ensembling (Rasmus et al., 2015; Tarvainen & Valpola, 2017; Laine & Aila, 2017) to address the stochastic variability in the global model and simply aggregates models based on reliable data amount. These strategies overlook the informative unlabeled clients with heterogeneous data, leading to suboptimal optimization. Therefore, a more comprehensive method for quantifying unlabeled clients is needed. In this work, we fill this gap by introducing SemiAnAgg, an anchor-based aggregation strategy in FedSemi, which harnesses the heterogeneity of unlabeled clients during valuation.
@@ 3 Preliminaries on Federated Semi-Supervised Learning
FedSemi Setting. Let us consider the generic Federated Semi-Supervised Learning (FedSemi) setting, which allows clients to be fully unlabeled, partially labeled, or fully labeled which is introduced in RSCFed (Liang et al., 2022) and CBAFed (Li et al., 2023). We denote the set of clients as $\left\{\mathcal{C}_{1},...,\mathcal{C}_{K}\right\}$, where each client $\mathcal{C}_{k}$ possesses a local dataset represented by ${\mathcal{D}_{client}}=\left\{\left\{{\left({X_{i}^{L},y_{i}^{L}}\right)}\right\}_{i=1}^{N^{L}},\left\{{\left({X_{i}^{U}}\right)}\right\}_{i=1}^{N^{U}}\right\}$. Here, $N^{L}$ and $N^{U}$ correspond to the number of labeled and unlabeled data instances, respectively. The objective is to derive a robust global model, $\theta_{glob}$, which effectively leverages all the available data across all clients. The most rigorous setting, as defined in (Liang et al., 2022; Li et al., 2023), is characterized by the majority of clients possessing fully unlabeled non-IID data, with $N^{L}=0$.
@@ 4.1 FedAvg-Semi: A Strong Aggregation Baseline for Fedsemi
${\theta_{global}}=\sum\limits_{k=1}^{K}{\hat{\lambda}_{1}\underbrace{\frac{{{N^{L}_{k}}}}{{{N^{L}_{total}}}}{\theta_{k}}}_{\text{Supervised}}}+\hat{\lambda}_{2}{\underbrace{\frac{{{\hat{N}^{U}_{k}}}}{{{\hat{N}^{U}_{total}}}}{\theta_{k}}}_{\text{Unsupervised}}},\;{N^{L}_{total}}=\sum\limits_{k=1}^{K}{{N^{L}_{k}}}\;\text{, and}\;{\hat{N}^{U}_{total}}=\sum\limits_{k=1}^{K}{{\hat{N}^{U}_{k}}},$ (2)
@@ 4.2 SemiAnAgg: Semi-Supervised Anchor-Based Aggregation
During the federated process, the client receives the averaged global model $\theta_{glob}$ from the server to update $\theta_{local}$. We can then extract $\hat{q}_{i}={\theta^{enc}_{glob}}(X_{i})$ along with the pseudo label $\hat{y}_{i}=\theta^{fc}_{glob}(\hat{q}_{i})$. Note that ${q_{i}}$ and $\hat{q}_{i}$ for each client are computed using the same $\theta^{enc}_{{glob}}$ and $\theta^{enc}_{{anch}}$, rendering client data, $\mathcal{D}_{client}$ as the only variable. We extract the features and compute the necessary pseudo-labels simultaneously, maintaining the computational complexity of $O(NE)$ per client per round, where $N$ represents the batch count and $E$ denotes the local epoch count.
@@
$\hat{r}_{k}={\sum\limits_{C}\hat{r}_{c}}\;\text{, and}\;{\theta_{global}}=\sum\limits_{k=1}^{K}{\hat{\lambda}_{1}\underbrace{\frac{{{N^{L}_{k}}}}{{{N^{L}_{total}}}}{\theta_{k}}}_{\text{Supervised}}}+{\hat{\lambda}_{2}\underbrace{\frac{\hat{r}_{k}}{\sum\limits\hat{r}_{k}}{\theta_{k}}}_{\text{Unsupervised}}},$ (4)
@@ 5.1 Experimental Setup
Baselines. In our evaluation, we compare our results with two state-of-the-art reproducible methods: RSCFed (Liang et al., 2022) and CBAFed (Li et al., 2023). To provide a more competitive baseline, we also present in Appendix D a study involving the SOTA in semi-supervised imbalanced learning under class distribution mismatch, implementing ACR (Wei & Gan, 2023) within the FedSemi framework.
@@ 5.2 Quantitative Comparisons with State-of-the-Art Methods
Results on the Balanced Global Setting. As shown in Table 1, SemiAnAgg achieves the best accuracy on two balanced datasets, SVHN and CIFAR-100, outperforming the compared methods in terms of accuracy. Specifically, SemiAnAgg surpasses RSCFed and CBAFed by 9.85%, and 0.12% on SVHN which is a relatively simple dataset. On a more challenging dataset, CIFAR-100, SemiAnAgg showcases a substantial improvement of 17.25% and 9.03%.
@@
Results with Partially Labeled Clients. In Table 3, we compare the generic FedSemi approaches in the scenario of having partially labeled clients on the SVHN and ISIC dataset, a subset of the setting. In this partially labeled setting, all clients possess 10% of their data as labeled, while the remaining 90% is considered unlabeled. Our SemiAnAgg method demonstrates higher performance over CBAFed (Li et al., 2023), achieving a 1.63% increase in accuracy on the SVHN dataset and a 4.41% improvement in “B-Acc” on ISIC. Further experiments and insights are provided in Appendix C
@@ 7 Limitations and Future Directions
In the absence of labeled data from unlabeled clients, SemiAnAgg approximates their contribution by comparing their data distribution through two consistently initialized models across clients. SemiAnAgg adds a layer of feature dictionary saving, nevertheless, it is negligible compared to the dataset size and the improvements obtained. Cosine similarity does not accurately represent the direction of the divergence given it shares only scalars. A more comprehensive approach that bounds the similarity within the unit sphere is necessary, nevertheless, it might raise privacy concerns. Despite its simple design, SemiAnAgg achieves state-of-the-art on four different benchmarks in FedSemi, compared to state-of-the-art federated semi-supervised learning (FedSemi). FedSemi is widely addressed in classification, whereas extending it to imbalanced regression (Yang et al., 2021; Wang & Wang, 2023) remains a challenging future task. SemiAnAgg’s lack of theoretical formulations is rather due to FedSemi complexity that has not been previously studied theoretically given the stochasticity of pseudo-labels highly dependent on initialization. We hope that our findings and insights inspire future approaches addressing the non-iid class mismatch and imbalanced problems in both centralized and decentralized settings.
@@
Appendix for “Semi-supervised Aggregation in Federated Learning”

### v2

# Learning Unlabeled Clients Divergence for Federated Semi-Supervised Learning via Anchor Model Aggregation

@@ 1 Introduction
Federated learning has emerged as a promising solution for learning in decentralized environments, where data centralization is often infeasible due to privacy concerns. Federated learning has gained considerable attention in machine learning tasks in natural image domains (McMahan et al., 2017) as well as medical image domains (Liu et al., 2021; Saha et al., 2023; Jiang et al., 2023). Due to the challenges of data and label heterogeneity, multiple methods such as MOON (Li et al., 2021a), FedDisco (Ye et al., 2023), FedFed (Yang et al., 2023), and FedRoD (Chen & Chao, 2022) have been developed, utilizing FedAvg (McMahan et al., 2017) as their baseline. While these methods have shown promise, they often assume that all clients have exhaustive and expert-level annotations, which is a requirement that is both time-consuming and labor-intensive. This renders them impractical in real-world cross-silo federated settings, such as those found in hospitals. For instance, in a federated learning framework where multiple hospitals collaborate to develop a shared model for complex tasks like lesion classification, some hospitals may provide fully labeled data for training, while others, with limited expert manpower, can only offer unlabeled or partially labeled data for model training. Therefore, developing methods that require minimal expert-level annotations in decentralized settings is a crucial area of research that deserves further attention.
Federated Semi-Supervised Learning (FedSemi) arises to alleviate exhaustive expert-level annotations by utilizing unlabeled data along with labeled data to benefit the global performance in many recent works (Zhang et al., 2023; Cho et al., 2023; Zhang et al., 2023; Wang et al., 2020; Jeong et al., 2021; Lin et al., 2021; Liu et al., 2021; Saha et al., 2023; Liang et al., 2022; Li et al., 2023; Kim et al., 2022; 2023b). Researchers have explored various settings for FedSemi, with the most generalizable setting considering fully unlabeled clients and other clients that can be either labeled or partially labeled, in either independent and identically distributed (IID) or non-IID settings (Liang et al., 2022; Li et al., 2023).
@@
Despite the effectiveness of these methods, existing approaches have two crucial limitations. First, their aggregation typically follows the standard FedAvg (McMahan et al., 2017), which aggregates clients based on the volume of local data, failing to account for the fact that labeled clients can generate more accurate information than unlabeled clients, regardless of data volume. For example, when the number of unlabeled clients is dominant, the global model becomes heavily influenced by the unlabeled clients, leading to limited performance primarily due to errors introduced by the pseudo-labels assigned to the unlabeled data. Second, existing methods treat the divergence of unlabeled clients, calculated in the gradient space, as noise due to the unreliability of the pseudo-labels used in training. It is worth mentioning that client divergence may reflect the presence of minority classes and unique attributes within the dataset, rather than mere noise.
@@
To this end, we propose a novel aggregation, SemiAnAgg, which learns the importance of unlabeled clients via a consistently randomly initialized model. Our key idea is that by feeding the local client data to the same global model and the same consistently initialized random anchor model, we can measure the importance of each unlabeled client. Specifically, our intuition is two-fold: 1) The optimal feature representation for client data should differ from a random representation, as random representations lack meaningful structure. Therefore, clients contributing to representations deviating from randomness are more likely to steer the model toward the global optimum. 2) Since all clients share the same global and anchor model, the inter-client data variability can be measured consistently. In other words, the expectation of the global model on client data that is more distant indicates a more diverse distribution covering unique information. Consequently, SemiAnAgg gives higher weights to these clients, harnessing their potential. Without requiring any direct sharing of data, features, or gradient information among the clients and maintaining the same communication costs as previous FedSemi approaches with minimal computation and storage overhead, SemiAnAgg achieves state-of-the-art results on four widely used benchmarks with improvements up to 9% in accuracy on CIFAR-100, 9.5% on the imbalanced CIFAR-100LT, and a 7.65% improvement in recall on the medical dataset ISIC-18.
@@ 2.2 Federated Semi-Supervised Learning
Federated Semi-Supervised Learning (FedSemi) addresses the decentralized learning of both unlabeled and labeled data while preserving privacy. FedSemi has been explored in three different settings. In the first two settings, labeled data is available on the server, as in FedMatch (Jeong et al., 2021), Semi-FL (Diao et al., 2022), and FedLID (Psaltis et al., 2023), or clients have partially labeled data, as seen in SemiFed (Lin et al., 2021) and FedLabel (Cho et al., 2023). The third setting, which we consider more realistic, involves fully unlabeled clients, while other clients can be either fully labeled or partially labeled. For the third setting, recent works have shifted from addressing clients with independent and identically distributed (IID) data, such as FedConsist (Wang et al., 2020) and FedIRM (Liu et al., 2021), to non-IID data, including RSCFed (Liang et al., 2022), IsoFed (Saha et al., 2023), and CBAFed (Li et al., 2023). This work focuses on the third setting, characterized by most clients with fully unlabeled and potentially non-IID data.
@@ 2.3 Federated Model Aggregation
Existing FedSemi (Liang et al., 2022; Li et al., 2023) methods regard non-robust clients or data as outliers and opt to minimize their influence. RSCFed (Liang et al., 2022) proposes to leverage unlabeled gradient divergence to eliminate non-robust (noisy) clients through average consensus inspired by RANSAC (Fischler & Bolles, 1981). However, RSCFed (Liang et al., 2022) does not consider cases when unlabeled clients are diverse due to unique information (i.e. underrepresented classes or attributes). IsoFed (Saha et al., 2023) proposes alternating model training between labeled and unlabeled clients in each round, which might result in significant catastrophic forgetting after each round. CBAFed (Li et al., 2023) introduced federated class adaptive pseudo labeling, which can be seen as curriculum-paced pseudo learning from a global perspective (Zhang et al., 2021). It uses hard temporal ensembling (Rasmus et al., 2015; Tarvainen & Valpola, 2017; Laine & Aila, 2017) to address the stochastic variability in the global model and simply aggregates models based on reliable data amount. These strategies overlook the informative unlabeled clients with heterogeneous data, leading to suboptimal optimization. Therefore, a more comprehensive method for quantifying unlabeled clients is needed. In this work, we fill this gap by introducing SemiAnAgg, an anchor-based aggregation strategy in FedSemi, which harnesses the heterogeneity of unlabeled clients during valuation.
@@ 3 Preliminaries on Federated Semi-Supervised Learning
FedSemi Setting. Let us consider the generic Federated Semi-Supervised Learning (FedSemi) setting, which allows clients to be fully unlabeled, partially labeled, or fully labeled which is introduced in RSCFed (Liang et al., 2022), IsoFed (Saha et al., 2023), and CBAFed (Li et al., 2023). We denote the set of clients as $\left\{\mathcal{C}_{1},...,\mathcal{C}_{K}\right\}$, where each client $\mathcal{C}_{k}$ possesses a local dataset represented by ${\mathcal{D}_{client}}=\left\{\left\{{\left({X_{i}^{L},y_{i}^{L}}\right)}\right\}_{i=1}^{N^{L}},\left\{{\left({X_{i}^{U}}\right)}\right\}_{i=1}^{N^{U}}\right\}$. Here, $N^{L}$ and $N^{U}$ correspond to the number of labeled and unlabeled data instances, respectively. The objective is to derive a robust global model, $\theta_{glob}$, which effectively leverages all the available data across all clients. The most rigorous setting, as defined in (Liang et al., 2022; Li et al., 2023), is characterized by the majority of clients possessing fully unlabeled non-IID data, with $N^{L}=0$.
@@ 4.1 FedAvg-Semi: A Strong Aggregation Baseline for Fedsemi
${\theta_{global}}=\sum\limits_{k=1}^{K}{\hat{\lambda}_{1}\underbrace{\left(\frac{{{N^{L}_{k}}}}{{{N^{L}_{total}}}}\right)}_{\text{Supervised}}{\theta_{k}}}+\hat{\lambda}_{2}{\underbrace{\left(\frac{{{\hat{N}^{U}_{k}}}}{{{\hat{N}^{U}_{total}}}}\right)}_{\text{Unsupervised}}{\theta_{k}}},\;{N^{L}_{total}}=\sum\limits_{k=1}^{K}{{N^{L}_{k}}}\;\text{, and}\;{\hat{N}^{U}_{total}}=\sum\limits_{k=1}^{K}{{\hat{N}^{U}_{k}}},$ (2)
@@ 4.2 SemiAnAgg: Semi-Supervised Anchor-Based Aggregation
During the federated process, the client receives the averaged global model $\theta_{glob}$ from the server to update $\theta_{local}$. From this, we can extract the feature representation $\hat{q}_{i}=\theta^{enc}_{glob}(X_{i})$, where $\theta^{enc}_{glob}$ refers to the global model’s encoder applied to the client data $X_{i}$. Additionally, the pseudo label is generated as $\hat{y}_{i}=\theta^{fc}_{glob}(\hat{q}_{i})$, where $\theta^{fc}_{glob}$ represents the global model’s fully connected layer used for classification. Note that ${q_{i}}$ and $\hat{q}_{i}$ for each client are computed using the same $\theta^{enc}_{{glob}}$ and $\theta^{enc}_{{anch}}$, rendering client data, $\mathcal{D}_{client}$ as the only variable. We extract the features and compute the necessary pseudo-labels simultaneously, maintaining the computational complexity of $O(BE)$ per client per round, where $B$ represents the batch count and $E$ denotes the local epoch count.
@@
$\hat{r}_{k}={\sum\limits_{C}\hat{r}_{c}}\;\text{, and}\;{\theta_{global}}=\sum\limits_{k=1}^{K}{\hat{\lambda}_{1}\underbrace{\left(\frac{{{N^{L}_{k}}}}{{{N^{L}_{total}}}}\right)}_{\text{Supervised}}{\theta_{k}}}+\hat{\lambda}_{2}{\underbrace{\left(\frac{\hat{r}_{k}}{\sum\limits\hat{r}_{k}}\right)}_{\text{Unsupervised}}{\theta_{k}}},$ (4)
@@ 5.1 Experimental Setup
Baselines. In our evaluation, we compare our results with state-of-the-art reproducible methods: RSCFed (Liang et al., 2022), IsoFed (Saha et al., 2023), and CBAFed (Li et al., 2023). To provide a more competitive baseline, we present in Appendix D and Appendix E additional baselines ablating local training strategies, specifically involving ACR (Wei & Gan, 2023) and SoftMatch (Chen et al., 2023) within the FedSemi framework.
@@ 5.2 Quantitative Comparisons with State-of-the-Art Methods
Results on the Balanced Global Setting. As shown in Table 1, SemiAnAgg achieves the best accuracy on two balanced datasets, SVHN and CIFAR-100, outperforming the compared methods in terms of accuracy. Specifically, SemiAnAgg surpasses RSCFed, IsoFed, and CBAFed by 9.85%, 9.21%, and 0.12%, respectively, on SVHN, which is a relatively simple dataset. On a more challenging dataset, CIFAR-100, SemiAnAgg showcases a substantial improvement of 17.25%, 16.74%, and 9.03% respectively.
@@
Results with Partially Labeled Clients. In Table 3, we compare the generic FedSemi approaches in the scenario of having partially labeled clients on the SVHN and ISIC dataset, a subset of the setting. In this partially labeled setting, all clients possess 10% of their data as labeled, while the remaining 90% is considered unlabeled. Our SemiAnAgg method demonstrates higher performance over CBAFed (Li et al., 2023), achieving a 1.63% increase in accuracy on the SVHN dataset and a 4.41% improvement in “B-Acc” on ISIC. Further experiments and insights are provided in Appendix C.
@@ 7 Limitations and Future Directions
In the absence of labeled data from unlabeled clients, SemiAnAgg approximates their contribution by comparing their data distribution through two consistently initialized models across clients. SemiAnAgg adds a layer of feature dictionary saving, nevertheless, it is negligible compared to the dataset size and the improvements obtained. Cosine similarity does not accurately represent the direction of the divergence given it shares only scalars. A more comprehensive approach, potentially involving multiple anchor models or feature representations, could better bound the similarity within the unit sphere. Nevertheless, sharing more independent features from clients might raise privacy concerns, which would need to be carefully addressed. Despite its simple design, SemiAnAgg achieves state-of-the-art on four different benchmarks in FedSemi, compared to state-of-the-art federated semi-supervised learning (FedSemi). FedSemi is widely addressed in classification, whereas extending it to imbalanced regression (Yang et al., 2021; Wang & Wang, 2023) remains a challenging future task. SemiAnAgg’s lack of theoretical formulations is rather due to FedSemi complexity that has not been previously studied theoretically given the stochasticity of pseudo-labels highly dependent on initialization. We hope that our findings and insights inspire future approaches addressing the non-iid class mismatch and imbalanced problems in both centralized and decentralized settings.
@@
Acknowledgments
This work was partially supported by grants from the National Natural Science Foundation of China (Grant No. 62306254) and the Hetao Shenzhen-Hong Kong Science and Technology Innovation Cooperation Zone (Project No. HZQB-KCZYB-2020083). Marawan Elbatel is supported by the Hong Kong PhD Fellowship Scheme (HKPFS) from the Hong Kong Research Grants Council (RGC), and by the Belt and Road Initiative from the HKSAR Government.
Appendix for “Semi-supervised Anchor-Based Aggregation in Federated Learning”
@@ Acknowledgments
    • In Appendix E, we provide an ablation study that examines the impact of different local client’s training strategies, including FlexMatch (Zhang et al., 2021) and SoftMatch (Chen et al., 2023).
@@ Appendix D Integration of Semi-Supervised Imbalance Learning Techniques
Appendix E Client Local Training Methods
In this section, we present an ablation study to evaluate the impact of different local training methods, namely FlexMatch (Zhang et al., 2021) and SoftMatch (Chen et al., 2023), on the ISIC-18 dataset.
Table 9: Ablation study on local training methods for FL. † indicates methods specifically tailored for FL.
Method Acc. (%) B-Acc (%)
FedAvg
CBAFed† (Li et al., 2023) 69.99 41.12
FlexMatch (Zhang et al., 2021) 62.16 31.73
SoftMatch (Chen et al., 2023) 66.99 38.02
FedAvg-Semi (our strong baseline)
FlexMatch (Zhang et al., 2021) 67.79 40.21
SoftMatch (Chen et al., 2023) 72.39 46.59
SemiAnAgg (ours)
FlexMatch (Zhang et al., 2021) 72.24 48.77
SoftMatch (Chen et al., 2023) 72.59 49.84
In Table 9, local semi-supervised learning methods show limited results due to the nature of FedAvg (McMahan et al., 2017), which does not fully account for the fact that labeled clients can provide more accurate information than unlabeled clients, regardless of data volume. While contributions such as CBAFed (Li et al., 2023) offer enhanced performance with a 9.41% increase in balanced accuracy, this still does not fully resolve the limitation.
Our FedAvg-Semi approach mitigates this limitation by balancing the empirical risk between labeled and unlabeled data, resulting in enhanced performance: an 8.48% increase in balanced accuracy for FlexMatch (Zhang et al., 2021) and an 8.57% increase in balanced accuracy for SoftMatch (Chen et al., 2023). Finally, by leveraging SemiAnAgg and considering the diversity of clients, we achieve the highest performance with a 7.65% increase in balanced accuracy for FlexMatch and 8.72% for SoftMatch compared with the state-of-the-art (SOTA) FedSemi approach, CBAFed (Li et al., 2023).